OpenStudy (anonymous):

Prove that lim as x-->3 (2x-6=0). (using the rigorous definition of limit)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3}2x-6=0\]

OpenStudy (anonymous):

you mean the epsilon-delta definition?

OpenStudy (anonymous):

I'm not sure. I just plugged in 3 for x to show that it equals 0 but I got it wrong.

myininaya (myininaya):

\[\text{ we want to show whenever } |x-a|<\delta \text{ then } |f(x)-L|< \epsilon \text{ if given } \\ \lim_{x \rightarrow a}f(x)=L\]

myininaya (myininaya):

so go ahead and plug in your L, f(x) and your a

myininaya (myininaya):

we are going to take |f(x)-L|<epsilon and try to solve for |x-a|

myininaya (myininaya):

this is not the proof this is just us trying to find a delta that works

myininaya (myininaya):

For example: \[\lim_{x \rightarrow 5}(2x+5)=15\] to prove this we must choose a delta such that whenever |x-a|<delta we then have |f(x)-L|<epsilon We want to show the then part eventually using the if part but first we must work backwards to find delta \[|f(x)-L|<\epsilon \\ |2x+5-15|<\epsilon \\ |2x-10|<\epsilon\] now remember we are trying to solve |f(x)-L|<epsilon for |x-a| so we can choose are delta you know that |x-a|<delta so anyways \[|2x-10|<\epsilon \\ |2| \cdot |x-5|<\epsilon \\ 2|x-5|<\epsilon \\ |x-5|< \frac{\epsilon}{2}\] so we will want to choose delta=epsilon/2 for our proof then for the actual proof part you will take this choosing of delta=epsilon/2 and show |f(x)-L|<epsilon